force F is equal to 3ti^ + 5j^ in acts on a body due to which its position varies as S is equal to 2t^2i^ - 5j^. Work done by this force in first 2 seconds is

Grade:11

1 Answers

Arun
6 years ago
Dear Vivek
 
Work done = F. dS
= (3t i + 5j).(2 t² i - 5j)
= 6t³ -25
 
At t= 2 sec
Work done = 6 *2³ -25 = 48- 25 = 23 Joules
 
Regards
Arun (askIITians forum expert)